Helsinki University of Technology Publications in Telecommunications Software and Multimedia

Teknillisen korkeakoulun tietoliikenneohjelmistojen ja multimedian julkaisuja Espoo 1999 TML-XX

IP network management

Liisa Uosukainen Tuomas Lilja Lasse Metso Seppo Ihalainen Jouni Karvo

Ossi Taipale

Helsinki University of Technology Department of Computer Science and Engineering Telecommunications Software and Multimedia Laboratory

Teknillinen korkeakoulu Tietotekniikan osasto Tietoliikenneohjelmistojen ja multimedian laboratorio CONTENTS

Contents 1

1 Introduction 2 1.1 The key sectors of network management . . . . . . . . . . . . . . . . . . . 2

2 Typical network management enviroments 3

2.1 OSI Network Management . . . . . . . . . . . . . . . . . . . . . . . . . . 3 OSI Management Model . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 OSI Information Model . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Communication Protocol for Transferring OSI Management Information . 3 OSI System Management Functions . . . . . . . . . . . . . . . . . . . . . 3

2.2 TMN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 TMN shortly . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 TMN sectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2.3 SS#7 Network Management . . . . . . . . . . . . . . . . . . . . . . . . . 4 Management Model . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 OMAP Applications Service Elements - ASE . . . . . . . . . . . . . . . . 4

2.4 TCP/IP Network Management . . . . . . . . . . . . . . . . . . . . . . . . 5 Concepts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Management Objects: SMI and MIB . . . . . . . . . . . . . . . . . . . . 5 Management Protocol: SNMP . . . . . . . . . . . . . . . . . . . . . . . . 5

2.5 Network Management Protocols . . . . . . . . . . . . . . . . . . . . . . . 5 SNMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 CMIP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

3 Future network management 7 3.1 SPIN’s Intelligent Network Management . . . . . . . . . . . . . . . . . . 7 3.2 Web-based network management . . . . . . . . . . . . . . . . . . . . . . 7 3.3 JMAPI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4 Telecommunications Management Processes 8 4.1 SMART TMN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

5 Network Management Technologies 9

5.1 System Management . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Opengroup - X/Open Systems Management . . . . . . . . . . . . . . . . . 9 System Management in Local Area Networks . . . . . . . . . . . . . . . . 9

5.2 CORBA Network Management . . . . . . . . . . . . . . . . . . . . . . . 9 CORBA Based Telecommunication Network Management System …. 9 Internetworking between CORBA and TMN Systems . . . . . . . . . . . 9

6 Commercial Network Management Products 10

Bibliography 12 1 INTRODUCTION

1.1 The key sectors of network management

The five key sectors of network management [6]:

• Fault Management enables the detection of a problem, fault isolation and correction to normal operation [6]. Other functions involved are alarm reporting, automatic recovery and different tests.

• Configuration Management is used to collect information from, send information to and control the operational states of network elements.

• Accounting is used to measure and establish charges for the use of the services provided.

• Performance Management offers functions to report and evaluate the operation of network and it’s elements. Statistical data is collected for the analysis and development of the network.

• Security Management is used to control the security features of the network. It includes user identification, different user levels and - TRAFFIC ENCRYPTING ???

2 TYPICAL NETWORK MANAGEMENT ENVIROMENTS

2.1 OSI Network Management

Computer networks usually contain a variety of equipment from several manufacturers. However, It should be possible to manage the whole network using only one network managing system. OSI Management Standards have been developed for that purpose. OSI Management encompasses the definition and implementation of management tools [3].

Base components of OSI network management are [3]

• Management Model

• Information Model

• communication Protocols for Transferring Management Information

• System Management Functions

OSI Management Model (SUORA LAINAUS a2:sta)

The management model is defined in terms of management applications that perform management activities in distributed manner by establishing associations between system management entities (agents and managers) [3]. The manager system manages the network resources, according to a defined policy, by issuing remote management requests to one or more agent processes. the agent process manages the real resources by executing the requests issued by the manager [3].

KUVA. MANAGERI-AGENTTI -KONSEPTI

OSI Information Model (SUORA LAINAUS a2:sta)

The OSI Information Model structures the management information according to a description of the resources to be managed in the network. Examples of suh resources may be a host, a routing table, a network device and an application process.The information model deals with managed objects that are abstractions of real resources for the purpose of management. They embody the management information for management applications [3].

Communication Protocol for Transferring OSI Management Information (SUORA LAINAUS a2:sta)

The transfer of the management information between agent and manager processes is performed using the CMIP (Common Management Information Protocol [3].

OSI System Management Functions OSI network management standard defines several system independent functions, that can be used by implementing managing applications that use the OSI application layer.

PITAISIKO SOVELLUSKERROKSESTA KERTOA ENEMMAN???

2.2 TMN

TMN shortly The management of a traditional telecommunication network is organized according to TMN (Telecommunication Management Network) standard. TMN:n consists of several ITU-T recommendations and it’s goal is to define the structure of the management network of the communications network. The purpose of this definition is to enable a connection between Network Elements and Operation Systems of different manufacturers’ devices.

TAULUKKO. ITU-T:n TMN SUOSITUKSIA

TMN is designed to support several issues concerning network management. Examples of such are maintenance, control, installation and designing the structures and services of the network.

KUVA. TMN:N SUHDE VERKKOON

TMN sectors ?????

• Monitoring

• Reporting

• Service Provisioning

• Network Develompment and Optimization

• Billing

2.3 SS#7 Network Management

SS#7 signalling protocol contains an Operations, Maintenance and Administration Part (OMAP), that offers a framework for operation and maintenance in SS#7 networks. OMAPlevel in SS#7 networks corresponds to the application layer in OSI networks.

SEURAAVA KAPPALE ON VAHAN HASSU: OMAP can be used with devices which offer SS#7 interface and have all the OMAP features implemented. Typically these are PDH and SDH devices, to which it might be difficult to implement other kind of management interface. However, OMAP is not being widely used....???

JATKUU....

Management Model SS#7 Management Model presents the dependence between different management components:

• System Management application Process (SMAP) monitores, controls and coordinates trough application layer protocols.

• Management Information Base (MIB) contains the information on performance and alarms collected by OMAP. MIB contains also configuration data that helps to manage the SS#7 network.

• Layer Management Interface (LMI) is not standardized due to it’s dependence on implementation. The data collected from protocol layer is stored to MIB using LMI.

• Layer Management Entity (LME) ????????????

• System Management Application Entity (SMAE) ??????????????

OMAP Applications Service Elements - ASE SS#7 OMAP defines two Applications Service Elements (ASE):

• Message Routing Verification Test (MRVT) ensures that the routing tables in the network are logical and that it does not routing cycles or other abnormalities.

• Circuit validation Test (CVT) ensures that two centers have equivalent data concerning the connections between them.

2.4 TCP/IP Network Management

Concepts The framework that is used to manage Internet-based networks is called “Internet-Standard Network Framework”. It differs from the corresponding OSI management framework????. Network management on TCP/IP-based networks is based on defining and solving real problems instead of al kinds of theorethical demands.

Management Concept in TCP/IP networks includes:

• Management Model:

TCP/IP Network Management System typically contains:

– Managed Nodes, devices that have network features. These devices include servers, workstations, printers, routers, etc.

– a Network Management Station (NMS). Management System should contain at least one Management Station. NMS is a server or a workstation that has the Network Management Protocol and management applications installed.

– a Network Management Protocol that is used to deliver requests between NMS and agents. NMS is able to read and write the variables of every Managed Node. The Network Management Protocol contains also Travelsar- and Trap-operations. Using Travelsar the NMS can find out what attributess do the Managed Nodes contain. Trap-operation is used by the Nodes to inform NMS about problems or unexpected events.

• Data Syntax:

The data used in network management is represented in the form of ASN.1 (Abstract Syntax Notation One). It defines the managed objects and the structure of the PDUs transferred by the Network Management Protocol.

ASN.1 is a machine-independent language that describes the structure of data. The data is encoded according to BER (Basic Encoding Rules).

Management Objects: SMI and MIB SMI (Structure of Management Information) defines several rules that can describe all the managed objects. The rules are independent of each other and the management protocol used.

MIB (Management Information Base) is a database that contains the management information of the network structure and the managed objects [4, pages 771–772]. MIB has a tree-like structure with the most detailed information stored in the leaves of the tree [4, pages 771–772].

Management Protocol: SNMP SNMP (Simple Network Managing Protocol) is widely used as the network managing protocol in TCP/IP-networks.

2.5 Network Management Protocols

SNMP SNMP (Simple Network Management Protocol) is the most widely used network management protocol. This is due to it’s simplicity, expandability, easy implementation and the fact that it poses only little stress on the network. It covers all the basic functions of network management [9].

SNMP is an asynchronous request/response protocol. This means that it does not have to wait for a response to a sent request before executing other tasks.

However, according to Lyn Vallillee, SNMP has some disadvantages: It contains large security gaps that can give network intruders access to the information carried along the network. Intruders could also potentially shut down some terminals. The biggest problem with SNMP though is that it is generally considered to be so simple that the information it deals with is neither detailed nor well-organized enough to deal with the expanding networks of the 1990’s. This is mainly due to the quick creation of SNMP, for it was never intended to lead network management into the 1990’s [9].

These problems were tried to be solved by establishing a new version of SNMP, called SNMPv2. This new version allows for more in-detail specification of variables, including the use of the table data structure for easier data retrieval. Also included are two new PDU’s that are used to manipulate the tabled objects. It has also added some security mechanisms that help combat the 3 largest security problems: privacy of data (to prevent intruders from gaining access to information carried along the network), authentication (to prevent intruders from sending false data across the network), and access control (which restricts access of particular variables to certain users, thus removing the possibility of a user accidentally crashing the network) [5].

However, SNMP still retains it’s disadvantages due to the fact that v2 never became a widely used protocol [9].

CMIP CMIP (Common Management Information Protocol) is a much more complicated and extensive network management protocol than SNMP. It improves on many of SNMP’s weaknesses, the security issues for instance, thus providing a more efficient network management enviroment [9].

Despite it’s advantages CMIP has not become a feasible choice instead of SNMP: Due to it’s complicity CMIP poses a lot of stress on the network and it’s implementation is very difficult [9].

3 FUTURE NETWORK MANAGEMENT

3.1 SPIN’s Intelligent Network Management

SPIN is a research project in the Institute for Information Technology at Canada’s National Research Council [2]. The SPIN Intelligent Network Management project studies and develops new agent-based technologies for controlling, planning and problem definition of heterogeneous networks. The application development of SPIN networks uses integration of the off-the-self NM components, and testing and usage of popular tools, such as HP openview.

3.2 Web-based network management

Doing network management operations using Internet/Intranet technologies is called Webbased Network Management. Web-based Network management comprises controlling network systems and / or data gathering, delivery of network management tasks and data analysis.

Basic applications of web based network management are web-based configuration and management of individual devices, advanced network wide management capabilities and web reporting of network status information.

For network element configuration with a WWW browser, a management agent with a HTML interface must be used. The HTML agent may configure the element using WWW forms and give reports as WWW pages.

Advanced network wide management capabilities seem to offer a WWW interface for traditional network management tools. Web reporting of network status means reporting statistics and query information of network elements on Intranet pages.

A pro for web based network management is the ability to use cheap hardware and software for user interfaces; the personnel may move physically and use the web interface for network management. However, it seems that web based network management mainly is a user interface improvement and does not add significantly to the actual network management.

3.3 JMAPI

Java Management API [7] is intended to provide a standard interface between different computers and network devices. The system can be used by Java programmes, and is created in alliance with Sun Micosystems, Bullsoft, Computer Associates, Exide Electronics, Jyra, Lumos Technologies, TIBCO, and Tivoli. Currently, only specification version 0.8 is available. Specification version 2.0 was intended for publication in March 1999.

4 TELECOMMUNICATIONS MANAGEMENT PROCESSES

4.1 SMART TMN

Smart TMN [8] is a program of Telemanagament Forum, a non-profit organization of dozens of product vendors and operators. The goal of SMART TMN is to present a larger scale model of telecommunications networks management.

The SMART TMN consists of four elements:

• Telecommunications Operations Map (TOM) which describes key business processes,

• Technology Integration Map, which contains recommendations of technologies to adopt for different management applications [1],

• Central Information Facility for information store for technical specifications, object models etc., and

• Catalys Projects are projects to validate technology concepts.

5 NETWORK MANAGEMENT TECHNOLOGIES

5.1 System Management

Opengroup - X/Open Systems Management System Management in Local Area Networks

5.2 CORBA Network Management

CORBA Based Telecommunication Network Management System Internetworking between CORBA and TMN Systems 6 COMMERCIAL NETWORK MANAGEMENT PRODUCTS

HP OpenView Hewlett-Packard HP OpenView has a wide variety of features. In addition to traditonal network management it includes system management and service management features. It provides users with Web/Java-based and traditional user interfaces. http://www.openview.hp.com

SUN Solstice Sun Microsystems Solstice is a traditional network management application. Different versions are provided for SOLARIS, Windows and Web/Java-based enviroments and for TMN and SNMP network management. http://www.sun.com/solstice

Java Dynamic Management Kit Sun Microsystems JDMK is the first application using JMAPI technology. It helps to develop Web-based management services using Java agents or JavaBeans for Management -tecnology. http://sun.com/software

TIVOLI IBM TIVOLI is a network management system that contains features that meet the requirements of especially demanding network enviroments. The product is aimed at both enterprises and network operators. http://www.tivoli.com

CYBERMANAGE Wipro Limited CYBERMANAGE is a Web-based network manager. It includes development tools. http://cybermanage.wipro.com/index1.html

SPECTRUM CABLETRON Systems http://www.cabletron.com/spectrum

INTEL LANDesk Intel http://www.intel.com/network/products/LANDesk srvr mgr.htm

ASANTEVIEW Asante Technologies, Inc. http://www.asante.com/products/p soft int.html

UNICENTER TNG Computer Associates http://www.cai.com

CONCORD Concord Communications http://www.concord.com/products.htm

CLEARSTATS RedPoint Network Systems http://www.redpt.com/ClearStats/ NETSCOUT WEBCAST Netscout http://www.netscout.com/Products/WebCast/body webcast.html BIBLIOGRAPHY

[1] SMART TMN^{T M} Technology Integration Map. Telemanagement Forum, Oct. 1998.

[2] SPIN homepage. http://www.iit.nrc.ca/SPIN public/english.html, 1999. Jun 2, 1999.

[3] Deri, L., and Mattei, E. An Object-Oriented Approach to the Implementation of OSI Management. Computer Networks and ISDN Systems 27, 9 (Aug. 1995), 1367–1385.

[4] Halsall, F. Data Communications, Computer Networks and Open Systems, 4 ed. Addison-Wesley, 1995.

[5] Stallings, W. SNMP, SNMPv2, and CMIP. Addison-Wesley, 1993.

[6] Stevenson, D. W. Network Management—What it is and what it isn’t. http://netman.cit.buffalo.edu/Doc/DStevenson, 1995. May 26, 1999.

[7] Sun Microsystems. Java^{T M} Management API (JMAPI).

http://www.javasoft.com/products/JavaManagement/, 1999. Jun 2, 1999.

[8] Telemanagement Forum. SMART TMN overview.

http://www.tmforum.org/pages/overview/tmfovr.html, 1999. Jun 2, 1999.

[9] Vallillee, L. SNMP and CMIP - An Introduction to Network Management. http://Home.InfoRamp.Net/ kjvallil/t/work.html, 1999. May 26, 1999.